### v1

# Enhancing Content-based Recommendation via Large Language Model

Abstract.
In real-world applications, users express different behaviors when they interact with different items, including implicit click/like interactions, and explicit comments/reviews interactions. Nevertheless, almost all recommender works are focused on how to describe user preferences by the implicit click/like interactions, to find the synergy of people. For the content-based explicit comments/reviews interactions, some works attempt to utilize them to mine the semantic knowledge to enhance recommender models. However, they still neglect the following two points: (1) The content semantic is a universal world knowledge; how do we extract the multi-aspect semantic information to empower different domains? (2) The user/item ID feature is a fundamental element for recommender models; how do we align the ID and content semantic feature space? In this paper, we propose a ‘plugin’ semantic knowledge transferring method LoID, which includes two major components: (1) LoRA-based large language model pretraining to extract multi-aspect semantic information; (2) ID-based contrastive objective to align their feature spaces. We conduct extensive experiments with SOTA baselines on real-world datasets, the detailed results demonstrating significant improvements of our method LoID.
Recommender System; Contrastive Learning; LLM;
Equal contributions to this work
Corresponding authors.
ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
1. Introduction
Background. With the boom of digital information, billions of user requests are produced daily, so recommender systems (RSs) have become an integral part of Internet platforms. To capture users’ interests more accurately, RSs have gone through several milestones, such as the logistic regression with hand-crafted features (e.g., FM (Rendle, 2010)), the neural networks (e.g., WideDeep (Cheng et al., 2016), YoutubeNet (Covington et al., 2016)), the sequential signal (e.g., DIN (Zhou et al., 2018), SIM (Pi et al., 2020)), and the multi-hop graph signal (e.g., PinSage (Ying et al., 2018), DGRec (Liangwei Yang, 2023)). In retrospect, these effective methods are based on the collaborative filtering (CF) idea and extend the boundaries of RSs. However, the CF framework also limits them under the case of cold-start and data sparsity problems. The reason is that the CF idea aims to mine the user/item pattern intelligence from data, to discover and recommend high-click candidate items for a user while it is hard to understand the users’ fine-grained and multi-aspect interests. In fact, instead of mining user preferences from massive user-item interaction logs, the user always leaves some reviews/comments to explain further his/her feelings about this interaction, which provides an explicit way to understand the users’ complex interests in language semantic space.
Related work. To extract valuable content semantic information, the pioneering works are formed as a rating prediction task: for a user-item pair in test set, give the historical user/item contents in training set (e.g., reviews), then predict their possible interaction rating. In early years, the DeepCoNN (Zheng et al., 2017) employed two convolutional neural networks (CNNs) towers to aggregate user/item content tokens individually to measure their dot-similarity, and the D-Attn (Seo et al., 2017) further extended DeepCoNN by introducing local and global attention mechanism to replace CNNs to aggregate tokens. Following the D-Attn, the ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018) focused on extracting the fine-grained multi-aspect semantic information and assigning different weights for aspects. The recent progress is RGCL (Shuai et al., 2022), which used BERT to generate the user-item content scores and then leveraged them as user-item graph edge weights to conduct a multi-hop graph neural network to make prediction.
Motivation. Although these methods raised model ability with content information, they ignore the following problems:
  • Transfer semantic knowledge across domains: Actually, RS needs to serve several domains simultaneously, such as electronics, clothing, books, etc. Since different domains always express different aspects of users’ interests (e.g., food is delicious, price is appropriate, etc.), the previous methods need to re-tune their semantic component for different domains, which is time-consuming.
  • Enhancing the correlation between content and ID: The user-item content and ID-interaction information can be seen as two different modalities (Li et al., 2023) connected by users. Nevertheless, previous methods focus on utilizing separate components to model the two corresponding content/ID spaces while ignoring how to exploit the correlation and align them in a unified space.
Our Work. To alleviate the above problems, we propose LoID, a LLM-based (Brown et al., 2020) model for transferring semantic knowledge across domains based on LoRA (Hu et al., 2021), and aligning content/ID information with contrastive objectives (Radford et al., 2021). It mainly includes two steps:
  1. (1)
    Pretraining semantic ‘plugin’: On the one hand, the ideal semantic information should act as a universal role to support all domains. On the other hand, different domains have their main aspects of semantic information that are related to recommendation. Considering such a trade-off, we borrow the LoRA strategy idea to train a small set of parameters for each domain, which could serve as plugins that can be added seamlessly to the target domain without further re-training.
  2. (2)
    Aligning the content/ID feature spaces: As discussed before, the content and ID information can be seen as two modalities of user-item interaction. To minimize their gap, we introduce the contrastive idea of maximizing content/ID features’ mutual information to align their feature spaces.
Finally, to validate our LoID effectiveness, we extensively test LoID under 11 different domain datasets to show its superior ability.
Contributions. Our contributions are summarized as follows:
  • We give a ‘plugin’ idea to transfer semantic knowledge, which sheds light on building a new paradigm for recommendation.
  • We devise a novel content/ID feature alignment objective.
  • We conducted detailed analyses with SOTA methods and LLMs.
Figure 1. Model Architecture of the proposed LoID.
2. Preliminary
2.1. Problem Statement
This work considers a brief task: For one domain dataset $\mathcal{D}$, it contains user-item ratings and textual contents, and each user-item interaction primarily comprises the following four elements: user ${\mathrm{u}}$, item ${\mathrm{i}}$, the corresponding rating $r_{\mathrm{u,i}}$, and the textual content token list $\textbf{t}_{\mathrm{u,i}}$ left by the user. Our model aims to predict ratings by historical content and user/item ID information.
Besides, to test the ‘plugin’ idea effectiveness, we further consider multi-domain scenarios, that is: given source domain content information $\{(\textbf{t}^{\mathrm{source_{1}}},r^{\mathrm{source_{1}}}),(\textbf{t}^{\mathrm{source_{2}}},r^{\mathrm{source_{2}}}),\dots\}$, predict target domain $\mathcal{D}_{\mathrm{target}}$ rating score.
2.2. Low-Rank Adaptation (LoRA)
Before going on, we first explain the LoRA (Hu et al., 2021), to show the basic idea of how to fast tuning an LLM (Devlin et al., 2018). Indeed, the unit block of LLM, transformer (Vaswani et al., 2017), consists of two parts: (Masked) attention and MLPs (FFM). The two parts introduce four pre-trained parameter matrices (e.g., $\mathbf{W}_{\mathrm{Q}},\mathbf{W}_{\mathrm{K}},\mathbf{W}_{\mathrm{V}},\mathbf{W}_{\mathrm{FFM}}$). Then, the challenge is: how to update those matrices without re-training them or involving new large parameters. Thereby, LoRA was proposed by assigning two small matrics for any pre-trained matrix $\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}\in\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}$:
(1) $\small\begin{split}\mathbf{W}^{\mathrm{LoRA}}=\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}+\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}},\end{split}$
where $\mathbf{W}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times d}$ is the modified weight matrix, $\mathbf{B}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{\mathrm{LoRA}}\in\mathbb{R}^{r\times d}$ are low-rank matrices$(r<<d)$. LoRA freezes the original large parameter matrix $\mathbf{W}$, introduces and updates it by $\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times d}$. Next, we can fine-tune the LLM by any supervised Task that without updating all parameters as follows:
(2) $\small\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{Task}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}),\quad\theta_{\mathrm{Task}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{Task}},\end{split}$
where $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ is extra added parameters and $\theta_{\mathrm{Task}}$ is all parameters [^1] In our work, we only introduce extra 4M $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ parameters to tune 110M $\theta^{\mathrm{pre}}_{\mathrm{BERT}}$ BERT..
3. Methodology
3.1. Overview
The architecture of our method LoID is illustrated in Figure 1, which includes two processes. In part (a), we first train the LoRA parameters of source domain as ”plug-ins” to enhance target domain prediction without further re-training. In part (b), we first extract historical contents of user/item to obtain user/item semantic representation, and then align the ID and semantic to make target domain rating prediction, note that the source domain LoRA plugin is an optional choice.
3.2. LoRA-based Encoder Pretraining
Source LoRA. To get users’ behavior data from the source domain, We leverage the LoRA strategy to pre-train the source domain [^2] For better understanding, we use BERT as LLM to show our method details.. In the pre-training task, we leverage rating prediction by $\mathbf{CLS}_{\mathrm{source}}$ token to predict rating directly:
(3) $\small\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{source}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{source}})&,\ \ \ \theta_{\mathrm{source}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{source}},\\ \mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}=\text{BERT}(\textbf{t}_{\mathrm{u,i}}^{\mathrm{source}},\theta_{\mathrm{source}}),&\quad\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}=\text{Predict}(\mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}),\\ \end{split}$
where BERT()($\cdot$)( ⋅ ) is the LLM forward procedure, Predict()($\cdot$)( ⋅ ) is an MLP to generate the prediction score $\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}$, then we adopt the Mean Squared Error (MSE) loss to optimize LoRA parameters:
(4) $\small\begin{split}\mathcal{L}_{\mathrm{pre-train}}^{\mathrm{source}}=\frac{1}{|\mathcal{D}_{\mathrm{source}}|}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{source}}}(\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}-r_{\mathrm{u,i}}^{\mathrm{source}})^{2},\end{split}$
where $|\mathcal{D}_{\mathrm{source}}|$ means the number of the samples of source domains, $\hat{r}_{\mathrm{u,i}}$/$r_{\mathrm{u,i}}$ represent predicted/real rating respectively. After pre-training all source domain’s LoRA parameters, we can use them as plugins to enhance BERT when dealing with other target domain tasks without further re-training.
3.3. Re-LoRA In Target Domain
Target LoRA. Enabling the model to adapt to the target domain, we introduce a LoRA module (i.e., Target LoRA) again in the target domain to fine-tune the target LoRA parameters. In the Re-LoRA process, we freeze the parameters of the BERT and source LoRA, and the only part that needs training is Target LoRA:
(5) $\footnotesize\begin{split}\mathbf{W}^{\mathrm{LoRA}}_{\mathrm{target}}&=\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}+\mathbf{B}_{\mathrm{source}}^{\mathrm{LoRA,Frozen}}\mathbf{A}_{\mathrm{source}}^{\mathrm{LoRA,Frozen}}+\mathbf{B}_{\mathrm{target}}^{\mathrm{LoRA}}\mathbf{A}_{\mathrm{target}}^{\mathrm{LoRA}},\\ \theta_{\mathrm{target}}=\{\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}&,\theta^{\mathrm{LoRA,Frozen}}_{\mathrm{source}}\}+\mathrm{LoRA(}\{\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\theta^{\mathrm{LoRA,Frozen}}_{\mathrm{source}}\},\mathrm{Task}_{\mathrm{target}}),\end{split}$
where $\theta_{\mathrm{target}}$ is the LLM part’s parameters of target domain. In the next, we explain how to align the language semantic and ID space.
3.4. ID-based Contrastive Learning
As the fundamental element of CF-based recommender work, the user/item IDs are indispensable in achieving personalized signals. How to align the space of semantic/ID is the key to making the content information more competitive for industrial RS.
User/Item Text Encoder. Considering the computation scale of LLM, we randomly extract historical ${\mathrm{k}}$ contents in the target domain ($\{\mathbf{t}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{t}_{\mathrm{u}}^{\mathrm{k}}\}$/$\{\mathbf{t}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{t}_{\mathrm{i}}^{\mathrm{k}}\}$) to describe the target domain user/item ${\mathrm{u}}/{\mathrm{i}}$ holistic preferences/properties:
(6) $\small\begin{split}\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{u}}^{\mathrm{k}},\theta_{\mathrm{target}}),\quad\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{i}}^{\mathrm{k}},\theta_{\mathrm{target}}),\end{split}$
where $\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}$ and $\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}$ are the embeddings for the user’s and item’s $k^{th}$ content.$\textbf{t}_{\mathrm{u}}^{\mathrm{k}}$ and $\textbf{t}_{\mathrm{i}}^{\mathrm{k}}$ are the corresponding textual content token list, and $\theta_{\mathrm{target}}$ stands for the LLM parameter on the target domain.
Attention Layer. On top of the user/item historical content $\mathbf{CLS}$ semantic information, we devise a novel attention mechanism to exchange the user/item semantic/ID information. Specifically, we integrate the semantic information $\{\mathbf{CLS}_{\mathrm{u}}^{1},\dots,\mathbf{CLS}_{\mathrm{u}}^{k}\}$ into the item representation $\mathbf{p}_{i}$, versus versa for user representation $\mathbf{p}_{u}$. By this mechanism, ID embedding and $\mathbf{CLS}$ token are fused, after which we obtain the updated item representation $\mathbf{v}_{i}$. Similarly, on the user side, we obtain ${\mathbf{v}_{u}}$ as follows:
(7) $\small\begin{split}\mathbf{v}_{i}&=\text{Attention}(\mathbf{p}_{i},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\}),\\ \mathbf{v}_{u}&=\text{Attention}(\mathbf{p}_{u},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\}),\\ \end{split}$
where $\mathbf{v}_{i}$ and $\mathbf{v}_{u}$ are the updated item and user representations.
Contrastive Loss. The Attention mechanism primarily focuses on different parts of the input sequence but does not explicitly consider the relationship between users and items. So, we conduct contrastive learning, enhancing the similarity between interactive users and items. For each updated representation $\mathbf{v}_{i}$/$\mathbf{v}_{u}$, treated as an anchor, we pair it with the original user/item representation as a positive sample $\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}$, and another representation in the batch as a negative sample $\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}$. The goal is to minimize the distance $\mathrm{D}(anchor,positive)$ and maximize $\mathrm{D}(anchor,negative)$. The loss function is minimized to ensure that the representations of similar items/users are closer in terms of the distance $\mathrm{D}$.
(8) $\small\begin{split}\mathrm{D}&=||anchor-sample||^{2}_{2},\\ \mathcal{L}_{\mathrm{cl}}=\max(&0,\Delta+\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{+}_{u})-\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{-}_{u}))\\ +\max(&0,\Delta+\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{+}_{i})-\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{-}_{i})),\\ \end{split}$
where ${\mathbf{v}_{i}}/\mathbf{v}_{u}$ is the representation of the updated item/user, ${\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}}$ represents the original feature representation of the user/item associated with the current content, as positive sample. ${\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}}$ represents the negative sample. ${\Delta}$ is a margin constant that enforces a ”safe distance” constraint on correctly classified samples, ensuring that the model’s classification results have sufficient robustness.
3.5. Model Optimization
Finally, we concatenate $\mathbf{v}_{u}$ and $\mathbf{v}_{i}$ to obtain the final score:
(9) $\small\begin{split}\hat{r}_{\mathrm{u,i}}=\text{Predict}(\text{Concat}(\mathbf{v}_{u},\mathbf{v}_{i})),\ \ \mathcal{L}_{\mathrm{rec}}=\frac{1}{{|\mathcal{D}_{\mathrm{target}}|}}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{target}}}(\hat{r}_{\mathrm{u,i}}-r_{\mathrm{u,i}})^{2},\end{split}$
where $|\mathcal{D}_{\mathrm{target}}$— means the number of the samples, $\hat{r}_{\mathrm{u,i}}$ and $r_{\mathrm{u,i}}$ represent predicted and real rating respectively.
Throughout the training process, the total loss comprises the MSE loss of rating prediction loss and the contrastive loss incurred during contrastive learning:
(10) $\small\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\lambda\mathcal{L}_{\mathrm{cl}},$
where $\mathcal{L}_{\mathrm{rec}}$ denotes the prediction loss, ${\lambda}$ is the weight assigned to the contrastive loss, and ${\mathcal{L}_{\mathrm{cl}}}$ represents the contrastive loss.
4. Experiments
4.1. Experimental Setup
Table 1. The experimental results on Amazon datasets (MSE, lower is better.)
Datasets DeepConn D-attn ALFM ANR BiGI RGCL LoID LoID(Elec) LoID(Elec &Movie)
Electronics 1.659 1.744 1.563 1.445 1.1433 0.9621 0.8662 $\star$ $\star$
Movies & TV 1.207 1.246 1.193 1.112 1.3489 0.9239 0.8353 0.8167 $\star$
CDs & Vinyl 0.980 1.014 0.956 0.914 1.1831 0.8180 0.6516 0.6295 0.6217
Amazon Instant Video 1.178 1.213 1.075 1.009 1.2437 0.9357 0.5177 0.4891 0.4864
Baby 1.442 1.507 1.359 1.258 1.5166 1.1414 0.7373 0.6911 0.6876
Digital Music 0.749 0.775 0.725 0.688 0.7731 0.7735 0.4696 0.4380 0.4365
Musical Instruments 1.160 1.224 1.072 1.034 1.3580 0.7211 0.5830 0.5179 0.5166
Office Products 1.569 1.650 1.474 1.337 1.6902 0.7001 0.6150 0.5909 0.5573
Patio, Lawn & Garden 1.622 1.696 1.510 1.403 1.7205 0.7049 0.6622 0.6358 0.6129
Pet Supplies 1.565 1.628 1.485 1.377 1.6185 1.2380 0.7685 0.7304 0.7249
Video Games 1.498 1.533 1.383 1.292 1.4450 1.0826 0.7822 0.7541 0.7409
Remark $\star$’ indicates that this domain is served as the source domain.
4.1.1. Datasets
We selected 11 categories  [^3] http://snap.stanford.edu/data/amazon/productGraph/categoryFiles/ of representative datasets from the Amazon dataset (Smith and Linden, 2017). Among them, the smallest dataset Musical Instruments includes 339,231 users, 83,046 items, 500,176 ratings, and the largest dataset Electronics comprises 4,201,696 users, 476,002 items, 7,824,482 ratings. Considering the different data-scale, we select three largest datasets (e.g., Electronics, Movies, and CDs) as our source domains, which contain relatively rich records of user-item contents and ratings. Following (Chin et al., 2018; Catherine and Cohen, 2017; Chen et al., 2018; Seo et al., 2017), we randomly partitioned our datasets into training, validation, and test sets with an 8:1:1 ratio.
4.1.2. Baselines
We compare LoID with several baselines, which can be categorized into three classes: (1) Single-aspect methods, such as DeepCoNN (Zheng et al., 2017) and D-Attn (Seo et al., 2017), extract features from the historical semantic information of users and items using CNN and attention mechanisms. (2) Multi-aspect methods, including ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018), aim to extract multiple semantic aspects and their respective importance. (3) GNNs-based methods, such as BiGI (Cao et al., 2021) and RGCL (Shuai et al., 2022), integrate adjacent node information from collaborative filtering.
4.1.3. Parameter Settings
In our method, the embedding size $d$ is fixed as 768, the dropout rate is fixed 0.5, the learning rate is set as 1e-5, the batch size is fixed as 4, and the training spans 5 epochs. In the LoRA module, the low-rank hyper-parameter $r$ selected from 16 to 48 with step length 8; the $\lambda$ selected from 0.2 to 0.5 with step length 0.1; the number of extracted user/item historical contents $k$ is is set as $[3\rightarrow 5\rightarrow 7\rightarrow 10]$. In the following section, we report LoID results under $r=32$, $\lambda=0.2$, and $k=5$ by default.
In the experiment, we report the DeepConn, D-attn, ALFM, and ANR results in the original literature (Chin et al., 2018) directly. Besides, in BiGI, the feature and hidden layer dimension are both set to 128, the dropout rate is fixed to 0.3, the step size is 0.001 and the number of GNN layers is set to 2. In RGCL, the number of GCN aggregation units and the number of output units of RGCL are both set to 64. The dropout rates of Edges, GCN, and nodes are set to 1.0, 0.7, and 0.3. Among our method and all baselines, we use the Adam (Kingma and Ba, 2015) algorithm to update parameters.
Table 2. Performance comparison of Domain Correlation.
Datasets Electronics Movies CDs
Sim MSE Sim MSE Sim MSE
Amazon Instant Video 0.07 0.489(+5%) 0.28 0.473(+9%) 0.16 0.485(+6%)
Digital Music 0.08 0.438(+6%) 0.15 0.436(+7%) 0.33 0.425(+9%)
Musical Instruments 0.22 0.518(+11%) 0.11 0.546(+6%) 0.21 0.540(+7%)
4.2. Performance Comparisons
Table 1 shows the performance of LoID on eleven datasets in terms of MSE metrics. Compared to other baselines, we note that aspect-aware recommendation methods such as ALFM and ANR consistently outperform DeepCoNN and D-Attn. We attribute this to the limitations of DeepCoNN and D-Attn, which lack a comprehensive model for the intricate decision-making process in user-item interactions. In addition, RGCL outperforms single-aspect and multi-aspect methods, indicating superior performance of graph neural networks. However, BiGI, despite incorporating graph networks, falls short compared to single-aspect methods. This suggests incorporating user textual semantic information is effective. We observe that LoID achieves statistically significant improvement over all SOTA baseline methods. This highlights that when using user contents, ID information should not be overlooked; instead, aligning them with semantic information in the same space can achieve positive performance. Moreover, LoID (Elec) results are superior to LoID, emphasizing enhanced performance with information from different domains. This demonstrates the effectiveness of extracting multi-aspect semantic information to empower various domains and our plugin idea effectiveness.
Multi-LoRA Merging. When we employ multiple LoRAs from different domains, the results of Multi-LoRA are superior to Single LoRA. We believe the reason is that multiple LoRAs bring richer and more extensive users’ data, more accurately describing historical behavior and preference tendencies.
4.3. Discussion of Domain Correlation Effect
To explore the semantic relationship of different domains, we conducted a more granular study of how to enhance further the transfer of semantic information to empower different domains. We consider whether the recommendation performance is correlated to the similarity between the target and source domains. First, we select 100 reviews from each dataset randomly, employ the Sentence-BERT (Reimers and Gurevych, 2019), and quantify the cosine similarity between datasets. Then, we assign different datasets as source domains to show the improvements compared with the origin LoID (as shown in Table 2). We can conclude that an increase in the similarity between domains leads to a corresponding rise in the performance improvement ratio. This implies that the transfer of our method becomes more effective when domains exhibit higher similarity.
4.4. Discussion of Different LLMs
Table 3. Performance comparison of Different LLMs.
Datasets BERT GPT2-medium
Precision Recall F1-score Precision Recall F1-score
Amazon Instant Video 0.7402 0.7578 0.7472 0.615 0.5999 0.606
Digital Music 0.7692 0.7767 0.7712 0.7244 0.7136 0.7183
Musical Instruments 0.7376 0.7436 0.7382 0.5534 0.5227 0.5332
This section investigates the impact of different LLMs in the pre-training phase, considering two distinct paradigms, i.e., GPT and BERT. Due to prompt-based GPT being unable to predict floating-point numbers, thus we formulate the prompt:
Input template: Give some example: {content1} is score {score1}, {content2} is score {score2}. Guess the score (The score should be between 1 and 5, where 1 means the lowest score, and 5 means the highest score) of {current content}, we think the score is?     Target template: {score}, {explanation}.
To ensure a fair comparison, we tuned them by LoRA and adopted PRF (Precision, Recall, and F1-score) as the evaluation protocol. Table 3 presents the results, indicating that the BERT outperforms GPT2-medium. This superiority may be attributed to BERT belonging to bi-directional encoder-framework LLM, which is more powerful under the regression task than decoder-framework LLM.
5. Conclusions
This paper introduces a simple yet effective approach named LoID, which includes two major components. (1) ‘Pre-training plugin’, we propose a flexible plugin framework that could transfer different domain semantic knowledge without re-training. (2) ‘Aligning semantic/ID space’, we devise a novel attention mechanism to connect the semantic and ID space, making our model easily applied in industrial RS. Extensive experiments reveal that LoID surpasses existing SOTA methods, and in-depth analyses underscore the effectiveness of our model components. In the future, we will explore the vision signal to improve our model ability.