### v2

# Enhancing Content-based Recommendation via Large Language Model

Abstract.
In real-world applications, users express different behaviors when they interact with different items, including implicit click/like interactions, and explicit comments/reviews interactions. Nevertheless, almost all recommender works are focused on how to describe user preferences by the implicit click/like interactions, to find the synergy of people. For the content-based explicit comments/reviews interactions, some works attempt to utilize them to mine the semantic knowledge to enhance recommender models. However, they still neglect the following two points: (1) The content semantic is a universal world knowledge; how do we extract the multi-aspect semantic information to empower different domains? (2) The user/item ID feature is a fundamental element for recommender models; how do we align the ID and content semantic feature space? In this paper, we propose a ‘plugin’ semantic knowledge transferring method LoID, which includes two major components: (1) LoRA-based large language model pretraining to extract multi-aspect semantic information; (2) ID-based contrastive objective to align their feature spaces. We conduct extensive experiments with SOTA baselines to demonstrate superiority of our method LoID.
Recommender System; Contrastive Learning; LLM;
Equal contributions to this work
Corresponding authors.
§The publication does not relate to the author’s work at Kuaishou.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679913isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
1. Introduction
Background. With the boom of digital information, billions of user requests are produced daily, so recommender systems (RSs) have become an integral part of Internet platforms. To capture users’ interests more accurately, RSs have gone through several milestones, such as the logistic regression with hand-crafted features (e.g., FM (Rendle, 2010)), the neural networks (e.g., WideDeep (Cheng et al., 2016), YoutubeNet (Covington et al., 2016)), the sequential signal (e.g., DIN (Zhou et al., 2018), SIM (Pi et al., 2020)), and the multi-hop graph signal (e.g., PinSage (Ying et al., 2018), DGRec (Liangwei Yang, 2023)). In retrospect, these effective methods are based on the collaborative filtering (CF) idea and extend the boundaries of RSs. However, the CF framework also limits them under the case of cold-start and data sparsity problems. The reason is that the CF idea aims to mine the user/item pattern intelligence from data, to discover and recommend high-click candidate items for a user while it is hard to understand the users’ fine-grained and multi-aspect interests. In fact, instead of mining user preferences from massive user-item interaction logs, the user always leaves some reviews/comments to explain further his/her feelings about this interaction, which provides an explicit way to understand the users’ complex interests in language semantic space. And the recent effort MoRec (Yuan et al., 2023) claimed that multi-modal information yield better results in sequential recommendation.
Related work. To extract valuable content semantic information, the pioneering works are formed as a rating prediction task: for a user-item pair in test set, give the historical user/item contents in training set (e.g., reviews), then predict their possible interaction rating. In early years, the DeepCoNN (Zheng et al., 2017) employed two convolutional neural networks (CNNs) towers to aggregate user/item content tokens individually to measure their dot-similarity, and the D-Attn (Seo et al., 2017) further extended DeepCoNN by introducing local and global attention mechanism to replace CNNs to aggregate tokens. Following the D-Attn, the ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018) focused on extracting the fine-grained multi-aspect semantic information and assigning different weights for aspects. The recent progress is RGCL (Shuai et al., 2022), which used BERT to generate the user-item content scores and then leveraged them as user-item graph edge weights to conduct a multi-hop graph neural network to make prediction.
Motivation. Although these methods raised model ability with content information, they ignore the following problems:
  • Transfer semantic knowledge across domains: Actually, RS needs to serve several domains simultaneously, such as electronics, clothing, books, etc. Since different domains always express different aspects of users’ interests (e.g., food is delicious, price is appropriate, etc.), the previous methods need to re-tune their semantic component for different domains, which is time-consuming.
  • Enhancing the correlation between content and ID: The user-item content and ID-interaction information can be seen as two different modalities (Li et al., 2023) connected by users. Nevertheless, previous methods focus on utilizing separate components to model the two corresponding content/ID spaces while ignoring how to exploit the correlation and align them in a unified space.
Our Work. To alleviate the above problems, we propose LoID, a LLM-based (Brown et al., 2020) model for transferring semantic knowledge across domains based on LoRA (Hu et al., 2021), and aligning content/ID information with contrastive objectives (Radford et al., 2021). It mainly includes two steps:
  1. (1)
    Pretraining semantic ‘plugin’: On the one hand, the ideal semantic information should act as a universal role to support all domains. On the other hand, different domains have their main aspects of semantic information that are related to recommendation. We borrow the LoRA strategy idea to train a small set of parameters for each domain, which could serve as plugins. Then we fused them by DARE  (Yu et al., 2023)without further re-training.
  2. (2)
    Aligning the content/ID feature spaces: As discussed before, the content and ID information can be seen as two modalities of user-item interaction. To minimize their gap, we introduce the contrastive idea of maximizing content/ID features’ mutual information to align their feature spaces.
Finally, to validate our LoID effectiveness, we extensively test LoID under 11 different domain datasets to show its superior ability.
Contributions. Our contributions are summarized as follows:
  • We give a ‘plugin’ idea to transfer semantic knowledge, which sheds light on building a new paradigm for recommendation [^1] Codes at https://github.com/cjx96/LoID..
  • We devise a novel content/ID feature alignment objective.
  • We conducted detailed analyses with SOTA methods and LLMs.
2. Preliminary
2.1. Problem Statement
This work considers a brief task: For one domain dataset $\mathcal{D}$, and each user-item interaction primarily comprises the following four elements: user ${\mathrm{u}}$, item ${\mathrm{i}}$, the corresponding rating $r_{\mathrm{u,i}}$, and the textual content token list $\textbf{t}_{\mathrm{u,i}}$ left by the user. Our model aims to predict ratings by historical content and user/item ID information.
Besides, to test the effectiveness of the ‘plugin’ idea , we further consider multi-domain scenarios, that is: given source domain content information $\{(\textbf{t}^{\mathrm{source_{1}}},r^{\mathrm{source_{1}}}),(\textbf{t}^{\mathrm{source_{2}}},r^{\mathrm{source_{2}}}),\dots\}$, predict target domain $\mathcal{D}_{\mathrm{target}}$ rating score.
2.2. Low-Rank Adaptation (LoRA)
Before going on, we first explain the LoRA (Hu et al., 2021), to show the basic idea of how to fast tuning an LLM (Devlin et al., 2018). Indeed, the unit block of LLM, transformer (Vaswani et al., 2017), consists of two parts: (Masked) attention and MLPs (FFN). The two parts introduce four pre-trained parameter matrices (e.g., $\mathbf{W}_{\mathrm{Q}},\mathbf{W}_{\mathrm{K}},\mathbf{W}_{\mathrm{V}},\mathbf{W}_{\mathrm{FFN}}$). Then, the challenge is: how to update those matrices without re-training them or involving new large parameters. Thereby, LoRA was proposed by assigning two small matrices for any pre-trained matrix $\mathbf{W}_{\mathrm{LLM}}^{\mathrm{Frozen}}\in\theta^{\mathrm{Frozen}}_{\mathrm{LLM}}$:
(1) $\footnotesize\begin{split}\mathbf{W}^{\mathrm{LoRA}}=\mathbf{W}_{\mathrm{LLM}}^{\mathrm{Frozen}}+\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}},\end{split}$
where $\mathbf{W}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times d}$ is the modified weight matrix, $\mathbf{B}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{\mathrm{LoRA}}\in\mathbb{R}^{r\times d}$ are low-rank matrices$(r<<d)$. LoRA freezes the original large parameter matrix $\mathbf{W}$, introduces and updates it by $\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}}\in\mathbb{R}^{d\times d}$. Next, we can fine-tune the LLM by any supervised Task that without updating all parameters as follows:
(2) $\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{Task}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{LLM}},\mathrm{Task}),\quad\theta_{\mathrm{Task}}=\theta^{\mathrm{Frozen}}_{\mathrm{LLM}}+\theta^{\mathrm{LoRA}}_{\mathrm{Task}},\end{split}$
where $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ is extra added parameters and $\theta_{\mathrm{Task}}$ is all parameters [^2] In our work, we only introduce extra 4M $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ parameters to tune 110M $\theta^{\mathrm{pre}}_{\mathrm{BERT}}$ BERT.. Among different domains, $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ has the same structure, so it is a proper semantic component for knowledge transfer  (Fu et al., 2024).
2.3. Drop and Rescale (DARE)
Assuming that we have fine-tuned the same LLM in different specific tasks and obtained several (supervised fine-tuned) SFT models as $\{\theta_{\mathrm{Task_{1}}},\theta_{\mathrm{Task_{2}}},\dots\}$, a challenge is: how to merge them as an omnipotent LLM? To answer it, DARE (Yu et al., 2023) was proposed, which devises a drop-then-rescale strategy on the gradient update sets $\{\delta_{\mathrm{Task_{1}}},\delta_{\mathrm{Task_{2}}},\cdots\}$ to make the multiple homogenous LLM merging, here we take the LoRA-Style SFT as an example:
(3) $\footnotesize\begin{split}\delta_{\mathrm{Task_{i}}}=\theta_{\mathrm{Task_{i}}}-\theta_{\mathrm{LLM}}=\mathbf{B}^{\mathrm{LoRA}}_{\mathrm{Task_{i}}}\mathbf{A}^{\mathrm{LoRA}}_{\mathrm{Task_{i}}}\\ \end{split}$
Next, DARE performs random drop on $\delta_{\mathrm{Task}}$ with probability $p$, and then rescales the remaining ones by a factor $1/(1-p)$.
(4) $\footnotesize\begin{split}\tilde{\delta}_{\mathrm{Task_{i}}}=\mathrm{Drop}(\delta_{\mathrm{Task_{i}}}),\quad\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{i}}}=\tilde{\delta}_{\mathrm{Task_{i}}}/(1-p)\end{split}$
where $p$ can reach 0.9 even 0.99, $\tilde{\delta}_{\mathrm{Task_{i}}}$ and $\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{i}}}$ represent the delta parameters after dropping and rescaling respectively.
(5) $\footnotesize\begin{split}\theta^{\mathrm{DARE}}_{\mathrm{LLM}}&=\mathrm{DARE(}\theta^{\mathrm{Frozen}}_{\mathrm{LLM}},\theta_{\mathrm{Task_{1}}},\theta_{\mathrm{Task_{2}}},\cdots,\theta_{\mathrm{Task_{n}}})\\ &=\theta_{\mathrm{LLM}}^{\mathrm{Frozen}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{1}}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{2}}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{3}}}+\cdots\end{split}$
where $\theta^{\mathrm{DARE}}_{\mathrm{LLM}}$ is the merged SFT LLM.
3. Methodology
3.1. Overview
The architecture of our method LoID is illustrated in Figure 1. In part (a), we train the LoRA parameters of source domains as ”plug-ins” to enhance target domain prediction without further re-training. In part (b), we extract historical contents of user/item to obtain user/item semantic representation, and then align the ID and semantic to make target domain rating prediction, note that the source domain LoRA plugin is an optional choice.
3.2. Source Domain LoRA Pretraining
Figure 1. Model Architecture of the proposed LoID.
Source LoRA. To get users’ behavior data from the source domain, We leverage the LoRA strategy to pre-train the source domain. In the pre-training task, we leverage rating prediction by $\mathbf{CLS}_{\mathrm{source}}$ token to predict rating directly, here we take BERT as default LLM:
$\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{source_{k}}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{source_{k}}})&,\ \ \ \theta_{\mathrm{source_{k}}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{source_{k}}},\\ \mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}=\text{BERT}(\textbf{t}_{\mathrm{u,i}}^{\mathrm{source_{k}}},\theta_{\mathrm{source_{k}}}),&\quad\hat{r}_{\mathrm{u,i}}^{\mathrm{source_{k}}}=\text{Predict}(\mathbf{CLS}^{\mathrm{source_{k}}}_{\mathrm{u,i}}),\\ \end{split}$
where $k$ denotes the $k^{th}$ domain in the source domains. where BERT()($\cdot$)( ⋅ ) is the LLM forward procedure, Predict()($\cdot$)( ⋅ ) is an MLP to generate the prediction score $\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}$, then we adopt the Mean Squared Error (MSE) loss to optimize LoRA parameters:
(6) $\footnotesize\begin{split}\mathcal{L}_{\mathrm{pre-train}}^{\mathrm{source}}=\frac{1}{|\mathcal{D}_{\mathrm{source}}|}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{source}}}(\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}-r_{\mathrm{u,i}}^{\mathrm{source}})^{2},\end{split}$
where $|\mathcal{D}_{\mathrm{source}}|$ means the number of the samples of source domains, $\hat{r}_{\mathrm{u,i}}$/$r_{\mathrm{u,i}}$ represent predicted/real rating respectively.
3.3. Target Domain Re-LoRA
After obtaining several source semantic ‘plugin’ LoRAs, we then transfer their general modality knowledge to target in two steps.
Source LoRA Merging. The first step is to merge these pre-trained source plugins with LLM, as discussed before, we follow DARE (Yu et al., 2023) idea to merge multiple LoRAs parameters.
(7) $\footnotesize\begin{split}\theta^{\mathrm{DARE}}_{\mathrm{BERT}}=\mathrm{DARE(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\theta_{\mathrm{source_{1}}},\theta_{\mathrm{source_{2}}},\cdots,\theta_{\mathrm{source_{n}}})\end{split}$
where $n\geq 0$ represents the avaliable source domains, note that the $n=0$ means ignoring all source domain LoRAs.
Target LoRA. Enabling the model to adapt to the target domain, we introduce a LoRA module (i.e., Target LoRA) again in the target domain. In the Re-LoRA, only the Target LoRA tuning its parameters; the merged BERT parameters are frozen:
$\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{target}}=\mathrm{LoRA(}\theta^{\mathrm{DARE,Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{target}})&,\ \ \ \theta_{\mathrm{target}}=\theta^{\mathrm{DARE,Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{target}},\\ \end{split}$
where $\theta_{\mathrm{target}}$ is the LLM part’s parameters of target domain.
3.4. ID-based Contrastive Learning
As the fundamental element of CF-based recommender work, the user/item IDs are indispensable in achieving personalized signals. How to align the independent feature space of semantic/ID is the key to making the content information more competitive for industrial RS. Therefore, we design an auxiliary task.
User/Item Text Encoder. Considering the computation scale of LLM, we randomly extract historical ${\mathrm{k}}$ contents in the target domain ($\{\mathbf{t}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{t}_{\mathrm{u}}^{\mathrm{k}}\}$/$\{\mathbf{t}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{t}_{\mathrm{i}}^{\mathrm{k}}\}$) to describe the target domain user/item ${\mathrm{u}}/{\mathrm{i}}$ holistic preferences/properties:
(8) $\footnotesize\begin{split}\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{u}}^{\mathrm{k}},\theta_{\mathrm{target}}),\quad\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{i}}^{\mathrm{k}},\theta_{\mathrm{target}}),\end{split}$
where $\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}$ and $\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}$ are the embeddings for the user’s and item’s $k^{th}$ content.$\textbf{t}_{\mathrm{u}}^{\mathrm{k}}$ and $\textbf{t}_{\mathrm{i}}^{\mathrm{k}}$ are the corresponding textual contents.
Attention Layer. On top of the user/item historical content $\mathbf{CLS}$ semantic information, we devise a novel attention mechanism to exchange the user/item semantic/ID information. Specifically, we integrate the semantic information $\{\mathbf{CLS}_{\mathrm{u}}^{1},\dots,\mathbf{CLS}_{\mathrm{u}}^{k}\}$ into the item representation $\mathbf{p}_{i}$, versus versa for user representation $\mathbf{p}_{u}$.
(9) $\footnotesize\begin{split}\mathbf{v}_{i}&=\text{Attention}(\mathbf{p}_{i},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\}),\\ \mathbf{v}_{u}&=\text{Attention}(\mathbf{p}_{u},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\}),\\ \end{split}$
where $\mathbf{v}_{i}$ and $\mathbf{v}_{u}$ are the updated item and user representations.
Contrastive Loss. The Attention mechanism primarily focuses on different parts of the input sequence, but it does not consider the relationship between users and items. So, we conduct contrastive learning, enhancing the similarity between interactive users and items. For each updated representation $\mathbf{v}_{i}$/$\mathbf{v}_{u}$, treated as an anchor, we pair it with the original user/item feature representation as a positive sample $\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}$, and another representation in the batch as a negative sample $\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}$. The goal is to minimize the distance $\mathrm{D}(anchor,positive)$ and maximize $\mathrm{D}(anchor,negative)$.
$\footnotesize\begin{split}\mathrm{D}&=||anchor-sample||^{2}_{2},\\ \mathcal{L}_{\mathrm{cl}}=\max(0,\Delta+\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{+}_{u})&-\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{-}_{u}))+\max(0,\Delta+\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{+}_{i})-\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{-}_{i})),\\ \end{split}$
where ${\mathbf{v}_{i}}/\mathbf{v}_{u}$ is the representation of the updated item/user, ${\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}}$ represents the positive sample, and ${\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}}$ represents the negative sample. ${\Delta}$ is a margin constant that enforces a ”safe distance” constraint on correctly classified samples.
3.5. Model Optimization
Finally, we concatenate $\mathbf{v}_{u}$ and $\mathbf{v}_{i}$ to predict the rating:
(10) $\footnotesize\begin{split}\hat{r}_{\mathrm{u,i}}=\text{Predict}(\text{Concat}(\mathbf{v}_{u},\mathbf{v}_{i})),\ \ \mathcal{L}_{\mathrm{rec}}=\frac{1}{{|\mathcal{D}_{\mathrm{target}}|}}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{target}}}(\hat{r}_{\mathrm{u,i}}-r_{\mathrm{u,i}})^{2},\end{split}$
where $|\mathcal{D}_{\mathrm{target}}$— means the number of the samples, $\hat{r}_{\mathrm{u,i}}$ and $r_{\mathrm{u,i}}$ represent predicted and real rating respectively.
Throughout the training process, the total loss comprises the rating prediction loss and the contrastive loss:
(11) $\footnotesize\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\lambda\mathcal{L}_{\mathrm{cl}},$
where $\mathcal{L}_{\mathrm{rec}}$ denotes the prediction loss, ${\lambda}$ is the weight assigned to the contrastive loss, and ${\mathcal{L}_{\mathrm{cl}}}$ represents the contrastive loss.
4. Experiments
4.1. Experimental Setup
Table 1. The experimental results on Amazon datasets (MSE, lower is better.)
Datasets DeepConn D-attn ALFM ANR BiGI RGCL LoID(w/o ID CL) LoID LoID(Elec) LoID(Elec &Movie)
Electronics 1.659 1.744 1.563 1.445 1.1433 0.9621 0.8738 0.8666 $\star$ $\star$
Movies & TV 1.207 1.246 1.193 1.112 1.3489 0.9239 0.8526 0.8362 0.8167 $\star$
CDs & Vinyl 0.980 1.014 0.956 0.914 1.1831 0.8180 0.7001 0.6558 0.6295 0.6170
Amazon Instant Video 1.178 1.213 1.075 1.009 1.2437 0.9357 0.5097 0.5094 0.4891 0.4731
Baby 1.442 1.507 1.359 1.258 1.5166 1.1414 0.7436 0.7275 0.6911 0.6888
Digital Music 0.749 0.775 0.725 0.688 0.7731 0.7735 0.4664 0.4709 0.4380 0.4195
Musical Instruments 1.160 1.224 1.072 1.034 1.3580 0.7211 0.5878 0.5816 0.5179 0.5322
Office Products 1.569 1.650 1.474 1.337 1.6902 0.7001 0.6422 0.6016 0.5909 0.5621
Patio, Lawn & Garden 1.622 1.696 1.510 1.403 1.7205 0.7049 0.6833 0.6703 0.6358 0.6281
Pet Supplies 1.565 1.628 1.485 1.377 1.6185 1.2380 0.7436 0.7679 0.7304 0.7196
Video Games 1.498 1.533 1.383 1.292 1.4450 1.0826 0.7994 0.7881 0.7541 0.7336
Remark $\star$’ indicates that this domain is served as the source domain. We report results for LoID and its variants under an average of 5 different random seed result settings.
4.1.1. Datasets
We selected 11 categories  [^3] http://snap.stanford.edu/data/amazon/productGraph/categoryFiles/ of representative datasets from the Amazon dataset (Smith and Linden, 2017). Among them, the smallest dataset Musical Instruments includes 339,231 users, 83,046 items, 500,176 ratings, and the largest dataset Electronics comprises 4,201,696 users, 476,002 items, 7,824,482 ratings. Considering the different data-scale, we select three largest datasets (e.g., Electronics, Movies, and CDs) as our source domains, which contain relatively rich records of user-item contents and ratings. Following (Chin et al., 2018; Catherine and Cohen, 2017; Chen et al., 2018; Seo et al., 2017), we randomly partitioned our datasets into training, validation, and test sets with an 8:1:1 ratio.
4.1.2. Baselines
We compare LoID with three classes of baselines: (1) Single-aspect methods, such as DeepCoNN (Zheng et al., 2017) and D-Attn (Seo et al., 2017), extract features from the historical semantic information of users and items using CNN and attention mechanisms. (2) Multi-aspect methods, including ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018), aim to extract multiple semantic aspects and their respective importance. (3) GNNs-based methods, such as BiGI (Cao et al., 2021) and RGCL (Shuai et al., 2022), integrate adjacent node information from collaborative filtering.
Table 2. Performance comparison of Domain Correlation.
Datasets Electronics Movies CDs
Sim MSE Sim MSE Sim MSE
Amazon Instant Video 0.07 0.489(+5%) 0.28 0.473(+9%) 0.16 0.485(+6%)
Digital Music 0.08 0.438(+6%) 0.15 0.436(+7%) 0.33 0.425(+9%)
Musical Instruments 0.22 0.518(+11%) 0.11 0.546(+6%) 0.21 0.540(+7%)
4.1.3. Parameter Settings
The embedding size $d$ ,dropout rate,learning rate and batch size is fixed as 768, 0.5, 1e-5 and 4. In the LoRA module, the low-rank hyper-parameter $r$ selected from 16 to 48 with step length 8; the $\lambda$ selected from 0.2 to 0.5 with step length 0.1; the number of extracted user/item historical contents $k$ is is set as $[3\rightarrow 5\rightarrow 7\rightarrow 10]$. Among all methods, we use the Adam (Kingma and Ba, 2015) to update parameters.
4.2. Performance Comparisons
Table 1 shows the performance of LoID on eleven datasets. All the experiments are repeated 5 times to sample historical contents in Section 3.4, and we report the (average) test MSE obtained when the validation MSE is the lowest. We observe that LoID achieves significant improvement over all SOTA baseline methods. We note that aspect-aware methods such as ALFM and ANR consistently outperform DeepCoNN and D-Attn. We attribute this to the fact that DeepCoNN and D-Attn lack a comprehensive model for the intricate decision-making process in user-item interactions. In addition, RGCL outperforms single-aspect and multi-aspect methods, indicating superior performance of graph neural networks. However, BiGI, though GNNs-based methods, falls short compared to single-aspect methods, without making use of reviews.
We observe that the results become worse when removing the contrastive learning module. Moreover, LoID (Elec) results are superior to LoID. This demonstrates the effectiveness of our LoRA-based plugin idea. When multiple LoRAs are employed, the results outperform when using single LoRA. We believe more LoRAs, more users’ behavior and preference data.
4.3. Discussion of Domain Correlation Effect
We investigate whether the performance is correlated to the similarity between the target and source domains. We select $n$ reviews from each dataset randomly, employ the Sentence-BERT (Reimers and Gurevych, 2019), to quantify the cosine similarity between the domains. In our experiments, we set $n$ to be 100. Then, we show the improvements compared with the original LoID (as shown in Table 2). We conclude that as the similarity between the source and target domains increases, there is a corresponding rise in the model’s performance.
4.4. Discussion of Different LLMs
Table 3. Performance comparison of Different LLMs.
Datasets BERT GPT2-medium
Precision Recall F1-score Precision Recall F1-score
Amazon Instant Video 0.7402 0.7578 0.7472 0.615 0.5999 0.606
Digital Music 0.7692 0.7767 0.7712 0.7244 0.7136 0.7183
Musical Instruments 0.7376 0.7436 0.7382 0.5534 0.5227 0.5332
We investigate the impact of different LLMs in the pre-training phase. We consider two distinct paradigms, i.e., GPT and BERT. Due to prompt-based GPT being unable to predict floating-point numbers, thus we formulate the prompt: Input template: Give some example: {content1} is score {score1}, {content2} is score {score2}. Guess the score (The score should be between 1 and 5, where 1 means the lowest score, and 5 means the highest score) of {current content}, we think the score is?     Target template: {score}, {explanation}. To ensure a fair comparison, we tuned them by LoRA and adopted PRF (Precision, Recall, and F1-score) as the evaluation protocol. Table 3 presents the results, indicating that the BERT outperforms GPT2-medium.
5. Conclusions
This paper introduces a simple yet effective approach named LoID, which includes two major components. (1) ‘Pre-training plugin’, we propose a flexible plugin framework that could transfer different domain semantic knowledge without re-training. (2) ‘Aligning semantic/ID space’, we devise a novel attention mechanism to connect the semantic and ID space, making our model easily applied in industrial RS. Extensive experiments reveal that LoID surpasses existing SOTA methods, and in-depth analyses underscore the effectiveness of our model components. In the future, we will explore the vision signal to improve our model ability.
Acknowledgement
This work is supported by the National Natural Science Foundation of China under Grant No. 62206128 and the Undergraduate Research Training Program of Nanjing University of Science and Technology (established in 2023) under Grant No. 2023066021B.